# Structure-aware Domain Knowledge Injection for Large Language Models

### v2

@@ 3.3 Structure-aware Supervised Fine-Tuning
In traditional knowledge injection paradigms, supervised fine-tuning serves as a tool to align the (continually) pre-trained models to interactive chatbots, e.g., through massive question-answering exercises (Cui et al., 2023; Qiu et al., 2024). However, most QA data augmentation or generation strategies aim at enlarging the quantity and enhancing the diversity of training samples (Xu et al., 2023a; Mukherjee et al., 2023; Liu et al., 2024b), which neglects the nature of the highly structured domain knowledge. Therefore, our structure-aware supervised fine-tuning (SSFT) technique focuses on eliciting models’ structured knowledge learned during the SCPT stage, adapting LLMs to interactive and reliable domain experts.
Fig. 2 (c) illustrates the idea of synthesizing SSFT samples guided by domain knowledge structures. Specifically, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap. For paths linking to a single knowledge point, we use the corresponding text content to prompt Llama3-70B (Dubey et al., 2024) to generate knowledge-intensive question-answering pairs. For paths with two or more branches, we prompt Llama3-70B with the knowledge path and textual contents to synthesize 2- or multi-hop QA samples, which require specific reasoning along the knowledge structure to derive from questions to answers. Fig. 5 presents several examples.
For every synthesized QA sample (${\bm{z}}$), we will prepend the relevant mindmap hierarchy to the answer, and add a CoT prompt in the question to construct another type of QA data (${\bm{z}}^{\prime}$) for SFT alignment. This design explicitly elicits the learned knowledge in models’ responses, teaching them how to apply the structured knowledge to address real-world problems. We use the two types of QA samples for training, as recommended by Qiu et al. (2024). During testing, we can either use the vanilla question as input to efficiently gather models’ answers, or take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.
Integrating with SCPT and SSFT, our StructTuning approach translates into remarkable efficacy and efficiency in domain knowledge injection, as comprehensively evaluated in the following sections.
@@ 4 Experiments
We design a comprehensive evaluation of our StructTuning through several experiments on two benchmarks. First, we investigate the free-form question-answering task on the LongBench (Bai et al., 2023) dataset, in order to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora). Then, we delve into the multi-choice question-answering task on MMedBench (Qiu et al., 2024), to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
4.1 Preliminary Investigation on Free-form Question-Answering
Table 1: Recall evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023) dataset. The best results are marked in bold, and the secondary results are marked with underlines. The backbone model is Llama2-7B (Touvron et al., 2023b).
@@ 4.1 Preliminary Investigation on Free-form Question-Answering
Datasets and Tasks. LongBench (Bai et al., 2023) is a multi-task benchmark tailored for open-book reading comprehension evaluation, where LLMs generate answers to given questions based on one or several input passages. To focus on knowledge injection, we turn the open-book evaluation into a closed-book QA task, where LLMs are trained on contextual passages and queried with questions only. We choose 7 subsets with 1,350 test examples from LongBench for single- and multi-document question-answering evaluation, and the remaining synthetic or code-orientated tasks are eliminated. More details are described in Sec. A.1.
Evaluation Metrics. To quantify the knowledge memorization degree, we mainly report the recall (Zhu & Li, 2023b) for models’ outputs against ground-truth answers for the free-form QA tasks. In Sec. B.3 we also evaluate our method with the F1-score measure for a thorough comparison.
Investigated Models. For LongBench, we mainly investigate the knowledge injection to Llama2-7B (Touvron et al., 2023b) to compare the open- and closed-book QA performance.
Implementation Details. For closed-book QA, the Llama2-7B model is continually pre-trained on 10,476 passages for 3 epochs using a batch size of 128. The initial learning rate is 2e-5, decayed to 0 at the end of training with a cosine scheduler. Then, we query Llama3-70B to generate 2,700 SSFT QA examples and remove those with over 0.5 F1-Score similarity to test samples to prevent data leakage. We train the knowledge-injected models for 1 epoch to avoid overfitting.
@@
Main Results. First, we use a CoT instruction of “Answer the question and explain why” to set a baseline for Open-Book QA (OBQA). As shown in Tab. 1, it presents a moderate knowledge comprehending capability in SingleDoc QA tasks (e.g., over 40% recall on Qasper and MFQA), and relatively poor knowledge coordination in MultiDoc QA subsets (e.g., around 10% recall on Musiq and Duzh) due to the attention drift on longer inputs.
Then, we try to inject passage content into LLMs with a conventional CPT+SFT paradigm to benchmark the Closed-Book QA (CBQA) baseline. Note that we use two types of question templates (refer to Sec. 3.3) during training, and use the CoT version during testing to elicit models’ memorized knowledge in their responses. However, Tab. 1 indicates such a knowledge injection approach is ineffective, as the knowledge recall is even lower than the OBQA baseline (i.e., 24.2% v.s. 26.9%).
On the contrary, our SCPT strategy has already achieved a considerable knowledge recall of 26.7%, which is higher than the CPT-SFT paradigm of 24.2% and approaching the OBQA baseline of 26.9%. It implies our structure-aware continually pre-trained model has successfully associated the relevant passages with their entire knowledge structure for the given question, which provides the knowledge path to seek targeted information to derive the answer.
Furthermore, our SSFT technique explicitly teaches the model to recall the learned knowledge and answer the questions, which further improves the knowledge recall to 31.5% and even surpasses the Open-Book QA setting. The results indicate the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions, as exemplified in Fig. 6.
Particularly, our method receives significant enhancements on MultiDoc QA subsets, which implies even though the total text content may exceed the models’ attention window and influence the OpenBook-QA, we can still chunk the content into several pieces and successfully inject the knowledge points into LLMs meanwhile preserving the whole knowledge structure.
Figure 1: MindMap Recall
F1-Score BERTScore
0.61 0.87
In addition, we also use lexical ROUGE-L (Lin, 2004) and semantic BERTScore (Zhang et al., 2020) to quantify the memorization of injected knowledge structures, by comparing the mindmap in models’ responses (as Fig. 6 displays) with ground-truth answers. The results in Fig. 1 indicate a relatively good memorization of the injected knowledge mindmap, emphasizing the efficacy of our SCPT strategy.
4.2 In-depth Evaluation for Multi-choice QA Application
Datasets and Tasks. MMedBench (Qiu et al., 2024) is a multilingual medical multi-choice QA benchmark, with 45,048 QA pairs for adapting LLMs to medical experts and 8,518 for testing. We collect 76M textbook corpora from MedTextBooks (Jin et al., 2020) and MMedC (Qiu et al., 2024) for medical knowledge injection and use the training/test split from MMedBench for SFT/evaluation. We also curate another two sizes of training sets (with 30M and 132M tokens) to validate our method’s scalability. Detailed setup is in Sec. A.2.
Evaluation Metrics. For multi-choice QA in MMedBench (Qiu et al., 2024), we follow the default setting to calculate the accuracy on six language subsets, as well as the averaged scores. Metrics are computed by lexical exact-matching on models’ responses, rather than maximum token probabilities.
Investigated Models. For MMedBench, we extend the investigated LLMs across model scales and architectures including Llama2-7B/13B (Touvron et al., 2023b), InternLM2-7B (Zheng et al., 2024), and the recent Llama3-8B (Dubey et al., 2024). We also compare our knowledge-injected models with other popular domain-specified LLMs, such as MedAlpaca (Han et al., 2023), ChatDoctor (Yunxiang et al., 2023), PMC-LLaMA (Wu et al., 2024), and MMedLM (Qiu et al., 2024) models.
Implementation Details. We first train LLMs for 3 epochs on medical textbooks with a batch size of 128. Then, we ask Llama3-70B to create structure-aware explanations for existing 45K QA samples in MMedBench’s training split and 33K extra entries by traversing extracted knowledge structures. Syntheses with overlapped options in the test set are removed. In the SFT phase, the learning rate is set as 1e-6 to avoid overfitting on such an amount of SFT samples, as suggested by Qiu et al. (2024).
Main Results. In Tab. 2, we present the overall performance across a series of LLMs on the testing split of MMedBench. The results demonstrate the promising enhancement achieved by our StructTuning technique, which translates into consistent improvements on the advanced InternLM2-7B (Zheng et al., 2024) and Llama3-8B (Dubey et al., 2024) models, and largely outperforms the previous domain-specific LLMs like PMC-LLaMA (Wu et al., 2024) and MedAlpaca (Han et al., 2023). Notably, our structure-aware knowledge injection approach, using merely 76M tokens curated from medical textbooks, achieves over 50% performance (4.46% v.s. 8.71%, 2.57% v.s. 4.96%) against the state-of-the-art MMedLM (Qiu et al., 2024) method, which is trained on a huge MMedC (Qiu et al., 2024) corpora of 25.5B tokens. Our approach shows much more efficiency in transforming pre-trained LLMs into domain experts by structured-aware knowledge injection, e.g., we only use 0.3% corpus to achieve comparable or even slightly better performance on English and Spanish subsets on top of Llama3-8B.
However, our method is currently unable to achieve 100% of knowledge injection efficacy against MMedLM, which may comprise two major reasons: (1) some knowledge in the training corpus has already been learned by the foundation model (e.g., a slight decrease in the Chinese subset), and (2) knowledge in the curated 76M training data cannot cover all testing scenarios (e.g., a considerable gap on the Japanese subset due to the absence of Japanese corpus, see Tab. A2). Therefore, we make a step to evaluate our approach’s scalability on various corpora sizes in Fig. 8.
@@ 4.2 In-depth Evaluation for Multi-choice QA Application
Table 2: Multiple-choice accuracy evaluation on MMedBench (Qiu et al., 2024). We report each model’s accuracy across six languages separately, with “Average” denoting the mean score over six languages. We also compare the data requirement (“#Token”) for medical knowledge injection.
@@
ChatDoctor 43.52 43.26 25.63 18.81 62.50 43.44 39.53 -
PMC-LLaMA 47.53 42.44 24.12 20.74 62.11 43.29 40.04 -
MedAlpaca 46.74 44.80 29.64 21.06 59.38 45.00 41.11 -
Llama2-7B 43.36 50.29 25.13 20.90 66.80 47.10 42.26 -
InternLM-7B 44.07 64.62 37.19 24.92 58.20 44.97 45.67 -
InternLM2-7B 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00 -
+MMed 61.74 80.01 61.81 52.09 80.47 67.65 67.30 +8.71 25.5B
+Ours 60.80 79.19 50.75 45.34 75.39 66.85 63.05 +4.46 76M
Llama3-8B 63.86 78.23 48.24 50.80 71.48 64.15 62.79 +0.00 -
+MMed 66.06 79.25 61.81 55.63 75.39 68.38 67.75 +4.96 25.5B
+Ours 66.77 77.44 53.27 51.61 74.61 68.49 65.36 +2.57 76M
Figure 8: Comparison of knowledge injection approaches.
@@
Approach’s Scalability. In Tab. 2, we use roughly 0.3% of the 25.5B tokens in MMedC Qiu et al. (2024) to evaluate knowledge injection methods. Here, another two training corpora sizes are curated: 30M and 132M, which take around 0.1% and 0.5% of 25.5B tokens respectively. We compare the vanilla CPT-SFT paradigm and our SCPT-SSFT strategy across those data settings The backbone LLM is InternLM2-7B (Zheng et al., 2024), and the training settings follow the main experiments. According to Fig. 8, our method consistently surpasses the vanilla paradigm by a large margin, emphasizing the efficacy and efficiency of domain knowledge injection. In particular, we can fit two performance-ratio scaling curves from the data points in Fig. 8 as:
$p_{v}\approx-0.04(\log r)^{2}+13.3\log r+100.0;\quad p_{s}\approx-1.11(\log r)^{2}+7.63\log r+133.0$ (3)
where $p_{v}$ and $p_{s}$ denote the relative performance enhancement (%) for vanilla and structure-aware knowledge injection, and $r$ refers to the corpus ratio. The scaling law indicates we can achieve comparable performance (100%) with only 5% of the total training corpus, significantly reducing the costs for LLMs’ domain adaptation. On the other hand, it also indicates our method may lead to 133% enhancement with a further 100% comprehensive data utilization. However, we have not empirically verified those predictions due to time and resource constraints, which we view as a temporal limitation of our current work and leave it to future investigations.
@@
Approach’s Generalization. In addition to the InternLM2 and Llama3 models in previous experiments, we also investigate the generalization ability of our SCPT+SSFT paradigm on Llama2 (Touvron et al., 2023b) model series. As shown in Tab. 3, our method leads to consistently significant improvements on 7B (+8.78%) and 13B (+6.17%) backbone models. The results further demonstrate the generalizability and scalability of our StructTuning strategy across model architectures and sizes.

### v3

@@ 3.3 Structure-aware Supervised Fine-Tuning
Traditional supervised fine-tuning aims to align the (continually) pre-trained models to interactive ChatBots through question-answering exercises (Cui et al., 2023; Qiu et al., 2024). Previous studies focus on enlarging the quantity and enhancing the diversity of training syntheses (Xu et al., 2023a; Mukherjee et al., 2023; Liu et al., 2024b) but neglect the highly structured domain knowledge. In contrast, our structure-aware supervised fine-tuning (SSFT) technique aims to elicit models’ structured knowledge learned during SCPT, adapting LLMs to interactive and reliable domain experts.
Fig. 2 (c) illustrates SSFT samples synthesis guided by domain knowledge structures. First, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap (the illustration of knowledge paths and branches is displayed in Fig. A2). For paths linking to a single knowledge point, we use the corresponding text content to prompt Llama3-70B (Dubey et al., 2024) to generate knowledge-intensive QA pairs. For paths with two or more branches, we prompt Llama3-70B with the knowledge path and textual contents to synthesize 2- or multi-hop QA samples, which require specific reasoning along the knowledge structure to derive from questions to answers. Fig. 4 presents several examples.
For every synthesized QA sample (${\bm{z}}$), we will prepend the relevant mindmap hierarchy to the answer, and add a CoT prompt in the question to construct another type of QA data (${\bm{z}}^{\prime}$) for SFT alignment. This design explicitly elicits the learned knowledge in models’ responses, teaching them how to apply the structured knowledge to address real-world problems. We use the two types of QA samples for training, as recommended by Qiu et al. (2024). During testing, we use the vanilla question as input to efficiently gather models’ answers to calculate accuracy, and take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.
Integrating with SCPT and SSFT, our StructTuning approach translates into remarkable efficacy and efficiency in domain knowledge injection, as comprehensively evaluated in subsequent sections.
@@ 4 Experiments
Table 1: Recall of Closed-Book QA (CBQA) on LongBench (Bai et al., 2023b). The best and secondary results are marked in bold and underlines, respectively. The base model is Llama2-7B.
@@
We design extensive evaluations of our StructTuning through several experiments on two benchmarks. First, we investigate the free-form question-answering task on the LongBench (Bai et al., 2023b) dataset, so as to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora). Then, we delve into the multi-choice question-answering task on MMedBench (Qiu et al., 2024), to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
4.1 Preliminary Free-form QA Investigation
Datasets and Tasks. Seven subsets with 1,350 test examples from LongBench (Bai et al., 2023b) are utilized to evaluate closed-book question-answering evaluation, where the answers can be directly found in corresponding passages. The 14K reading-comprehension passages are used for knowledge injection via CPT/SCPT, and another 2,700 QA samples are generated for SFT/SSFT. Details are described in Sec. A.1.
Evaluation Metrics. Here we first report the recall (Zhu and Li, 2023b) for models’ outputs against ground-truth answers to quantify the knowledge memorization degree. In Sec. B.9 we also evaluate F1-score for a thorough comparison.
Investigated Models. We mainly investigate Llama2-7B (Touvron et al., 2023b) to compare the knowledge injection performance.
Implementation Details. We train all models on LongBench passages for 3 epochs using a batch size of 128, and train for 1 epoch on synthetic SFT data to avoid overfitting. The learning rate is 2e-5.
Main Results. We first try to inject passage content into LLMs with a conventional CPT+SFT paradigm for the baseline, and use CoT instructions during testing to elicit models’ memorized knowledge in their responses. However, Tab. 1 indicates such an injection approach is ineffective, as the knowledge recall is only 24.2%.
On the other hand, our SCPT strategy achieved a higher knowledge recall of 26.7%. It implies the model has successfully associated the relevant passages with their entire knowledge structure for the given question, especially on multi-doc QAs that require more complex information retrieval and reasoning on multiple passages. Furthermore, our SSFT technique continues to improve the knowledge recall to 31.5%, which indicates the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions. Fig. 5 showcases a representative example.
@@ 4.1 Preliminary Free-form QA Investigation
In Sec. B.8, we also validate relatively good memorization of injected knowledge structures, further emphasizing our SCPT’s efficacy.
@@
Table 2: Multiple-choice evaluation on MMedBench (Qiu et al., 2024). We report separate accuracies across six languages, with “Average” denoting the mean score. “#Token” denotes required data for knowledge injection.
@@
ChatDoctor (Yunxiang et al., 2023) 43.52 43.26 25.63 18.81 62.50 43.44 39.53 -
PMC-LLaMA Wu et al. (2024) 47.53 42.44 24.12 20.74 62.11 43.29 40.04 -
MedAlpaca Han et al. (2023) 46.74 44.80 29.64 21.06 59.38 45.00 41.11 -
Llama2-7B Touvron et al. (2023b) 43.36 50.29 25.13 20.90 66.80 47.10 42.26 -
InternLM2-7B Zheng et al. (2024) 57.27 77.55 47.74 41.00 68.36 59.59 58.59 -
Llama3-8B (AI, 2024) 63.86 78.23 48.24 50.80 71.48 64.15 62.79 +0.00 -
Llama3+MMed Qiu et al. (2024) 66.06 79.25 61.81 55.63 75.39 68.38 67.75 +4.96 25.5B
Llama3+StructTuning (Ours) 66.77 77.44 53.27 51.61 74.61 68.49 65.36 +2.57 76M
Llama3+StructTuning (Ours) 65.36 79.04 56.28 55.47 80.47 69.80 67.74 +4.95 1.2B
4.2 In-depth Multi-choice QA Evaluation
Datasets and Tasks. We take several corpus sizes from MMedC (Qiu et al., 2024) for CPT/SCPT, with 45K QA data from MMedBench (Qiu et al., 2024)’s training set for SFT/SSFT. Models are evaluated on six multi-choice subsets from MMedBench, where LLMs should make real-world patient diagnoses with adequate reasoning on medical knowledge. Detailed setup is in Sec. A.2.
Evaluation Metrics. We follow the default setting to calculate the accuracy on six language subsets and the averaged scores. Metrics are computed by lexical exact-matching on models’ responses, rather than maximum token probabilities.
Investigated Models. We extend the investigation across model architectures and scales, including Llama2-7B/13B (Touvron et al., 2023b), InternLM2-7B (Zheng et al., 2024), and Llama3-8B (Dubey et al., 2024). Other popular medical LLMs (Han et al., 2023; Wu et al., 2024; Qiu et al., 2024) are also included for a thorough comparison.
Implementation Details. Following Qiu et al. (2024), models are first trained for 3 epochs on medical corpora with a learning rate of 2e-5, and then fine-tuned for 1 epoch to avoid overfitting. The detailed setup is displayed in Sec. A.2.
Figure 6: Knowledge injection’s scalability.
Main Results. The results in Tab. 2 demonstrate the promising enhancement achieved by our StructTuning technique largely outperforming the previous domain-specific LLMs like PMC-LLaMA (Wu et al., 2024) and MedAlpaca (Han et al., 2023). Notably, our structure-aware knowledge injection approach, using merely 76M tokens (0.3%) curated from medical textbooks, achieves over 50% performance (2.57% v.s. 4.96%) against the state-of-the-art MMedLM (Qiu et al., 2024) method, which is trained on the entire MMedC (Qiu et al., 2024) corpora of 25.5B tokens. After scaling up the training tokens to 1.2B (around 5%), our method makes nearly 100% improvements to the average accuracy, significantly reducing the training cost of traditional knowledge injection approaches.
@@ 4.2 In-depth Multi-choice QA Evaluation
Approach’s Scalability. We further curate a series of training corpora sizes to investigate our method’s scalability in-depth: 30M, 76M, 132M, 250M, and 1.2B, which respectively take around 0.1%, 0.3%, 0.5%, 1%, and 5% of 25.5B tokens. The vanilla CPT-SFT paradigm and our SCPT-SSFT strategy are thoroughly compared across those data settings. According to Fig. 6, our method consistently surpasses the vanilla paradigm by a large margin, emphasizing the efficacy and efficiency of domain knowledge injection. In particular, we fit two performance-ratio scaling curves from the data points in Fig. 6 as:
$p_{v}\approx-0.04(\log r)^{2}+13.3\log r+100.0$ (3)
$p_{s}\approx-1.11(\log r)^{2}+7.63\log r+133.0$
where $p_{v}$ and $p_{s}$ denote the relative performance enhancement (%) for vanilla and structure-aware knowledge injection, and $r$ is the corpus ratio.
In fact, we use the scaling law derived from 0.1%, 0.3%, 0.5%, and 1% points to predict that achieving 100% performance would require 5% of the data corpus, which has been confirmed in Tab. 2. On the other hand, it also indicates our method may lead to 133% enhancement with a further 100% comprehensive data utilization, further validating the effectiveness and scalability of our method.
Table 3: Generalization to various model architectures and sizes. Here we use 76M training corpus.
Model Size Adaptation Accuracy
InternLM2 7B CPT+SFT 58.59
SCPT+SSFT 63.05
Llama2 7B CPT+SFT 42.26
SCPT+SSFT 51.04
Llama2 13B CPT+SFT 48.33
SCPT+SSFT 54.50
Adaptation English Chinese Japanese French Russian Spanish Average
SFT 44.54 32.81 26.63 15.27 53.91 42.30 35.91
CPT + SFT 46.27 32.57 26.13 17.36 50.00 40.63 35.49
SCPT + SFT 46.50 32.14 20.10 18.17 53.91 39.97 35.13
SCPT + SSFT 49.96 32.63 22.11 17.52 51.17 41.28 35.78
SCPT + SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
RAG 38.12 29.22 22.61 23.34 53.91 36.47 33.95
AdaptLLM (Cheng et al., 2023) 46.79 33.80 20.60 14.15 53.12 42.34 35.03
RAFT (Zhang et al., 2024) 43.60 32.34 21.11 14.95 50.39 42.16 34.09
@@
Approach’s Generalization. In Tab. 3, we also validated the performance on Llama2 (Touvron et al., 2023b) and InternLM2 (Zheng et al., 2024) model series by using 76M tokens for knowledge injection. Our method leads to consistently significant improvements on InternLM2-7B (+4.46%), Llama2-7B (+8.78%) and Llama2-13B (+6.17%) backbone models, further demonstrating the generalizability and scalability of our StructTuning across model architectures and sizes. Detailed results are presented in Tab. A6.
Ablation Studies. We perform a comprehensive ablation study on MMedBench’s English subset in Tab. 4. As suggested by Qiu et al. (2024), we use the English textbooks (Jin et al., 2020) (26M tokens) to compare vanilla and structure-aware CPT, paired with corresponding SFT strategies. In particular, “SFT” uses vanilla SFT with 10K QA samples from MMedBench’s training split, while “SSFT” applies structure-aware SFT on the same questions, enhancing answers with Llama3-70B-generated knowledge explanations (Sec. 3.3). “SSFT*” further includes 8K additional structure-aware QA pairs, totaling 18K training entries. Training hyperparameters align with the main experiment.
In Tab. 4, the CPT+SFT paradigm improves accuracy by 1.73%, while SCPT with vanilla SFT achieves a higher 46.50%. Combining SCPT with SSFT boosts performance significantly (49.96% v.s. 44.54%), highlighting the importance of structured knowledge elicitation. Adding 8K extra QA pairs (“SSFT*”) further improves performance across five subsets, demonstrating a surprising cross-lingual knowledge transfer (Lai et al., 2023; Qin et al., 2024). After SSFT, LLMs effectively use knowledge injected in one language to solve problems in others, surpassing traditional SFT. Additional comparisons in Sec. B.6 confirm that structure-aware syntheses can enhance knowledge application better than random syntheses.
In addition, we observe that the commonly used RAG (Lewis et al., 2020) strategy does not bring significant advantages to the MMedBench evaluation. The main reason lies in the gap between the pre-training corpus (comprising official knowledge statements from textbooks) and evaluated QA samples (originating from practical diagnosis records). Knowledge injection by (S)CPT and (S)SFT shows more advantages in this situation. In-depth investigations can be found in Sec. B.7.
Comparision with Other Methods. We also compare two advanced knowledge injection methods in Tab. 4 to further demonstrate our StructTuning’s efficacy: (1) AdaptLLM (Cheng et al., 2023): domain knowledge injection by appending reading comprehension QAs to each CPT chunk, and (2) RAFT (Zhang et al., 2024): improving LLM’s robustness to domain-specific retrieval-augmented generation using noisy retrieval-augmented SFT samples. According to the experimental results, AdaptLLM (Cheng et al., 2023) brings negligible improvement in the final performance (e.g., 46.79% v.s. 46.27% on the English subset), indicating such a chunk-level reading comprehension augmentation during CPT cannot help LLMs capture the entire structured domain knowledge. Concurrently, RAFT (Zhang et al., 2024) causes even worse performance, since the retrieval process introduces too many unrelated chunks and hurts LLM’s QA judgments, especially when there exists a significant gap between user query and knowledge chunks in the medical diagnosis scenario.
5 Conclusion
This work pioneers in incorporating structure-aware methodologies to enhance domain knowledge injection into large language models. Through a novel SCPT-SSFT paradigm, we have set a new precedent for adapting LLMs to specialized domains, and the promising and scalable results underscore the viability and potential of our method. We hope to inspire further research in efficient and effective domain adaptation in the LLM community, moving a step closer to models that can truly emulate human intelligence.
Limitation. Our two-stage strategy introduces added computational complexity, where taxonomy extraction and data reorganization are required in the SCPT phase, and extra QA syntheses are optionally applied in the SSFT stage. In Appendix B, we provide further discussion with extensive empirical experiments. Despite the additional computational overhead introduced, our method achieves greater overall benefits and can reduce the reliance on large-scale LLMs (e.g., 70B models). We will delve into the investigations in future work.
Appendix A Implementation Details
A.1 Detailed Setup on LongBench
Dataset Composition. To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench (Bai et al., 2023b) across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
  • Single-Doc QA. For single-document QA, we take three subsets from LongBench: (1) Qasper (Dasigi et al., 2021), featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA (Bai et al., 2023b), manually curated from multiple data sources and annotated by Ph.D. students; and (3) MultiFieldQA-zh, the Chinese split also provided by Bai et al. (2023b), covering multiple Chinese scenarios. MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others subsets include 200 pieces of test samples respectively.
  • Multi-Doc QA. Multi-document QA requires LLMs to extract and combine information from multiple documents to derive the answer, which is generally more challenging than single-doc QA. We take four multi-hop QA datasets: (1) HotpotQA (Yang et al., 2018), containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA (Ho et al., 2020), involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue (Trivedi et al., 2022), carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader (He et al., 2017), developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al. (2023b) to reduce the data noise. Each subset has 200 test samples.
In Single-Doc QA, we extract knowledge structures for each single passage; in Multi-Doc QA, we identify the knowledge structure across multiple passages for each test sample. There are ultimate 1350 question-answer-passage(s)-(knowledge)structure quadruples to evaluate knowledge injection approaches on LongBench.
SFT Data-Synthesis. We query Llama3-70B to generate 2,700 QA examples and remove those with over 0.5 F1-Score similarity to test samples to prevent data leakage. During inference, when the model can generate correct answers (corresponding to specific knowledge points) that haven’t been seen during the SFT stage, we can ensure the knowledge is injected at the CPT stage and SFT only enhances the instruction-following capability. In practice, merely 13 out of 2700 (around 0.5%) synthetic data have over 0.5 F1-Score and are thus filtered out from the SFT data.
Tab. A1 statistics the semantic similarity (measured by BERTScore (Zhang et al., 2020)) between generated and GT questions and answers, and the results emphasize there is no knowledge leakage in the generated SFT data (they share poor semantic similarity across questions, answers, and QAs).
Table A1: Similarity statistics on synthetic SFT data and LongBench’s test samples.
Target Question Answer Question-Answer
BERTScore 0.277 0.106 0.093
Table A2: Sample statistics on MMedBench’s QA data.
Split English Chinese Japanese French Russian Spanish Total
Train 10,178 27,400 1,590 2,171 1,052 2,657 45,048
Test 1,273 3,426 199 622 256 2,742 8,518
Table A3: Sample statistics on different training data settings.
Stage Ratio English Chinese Japanese French Russian Spanish Total
CPT 0.1% 6.9M 8.8M - 4.1M 4.9M 5.4M 30.1M
CPT 0.3% 26.1M 21.5M - 8.1M 10.3M 10.1M 76.1M
CPT 0.5% 35.9M 27.2M 4.5M 14.0M 24.9M 26.1M 132.6M
CPT 1.0% 44.1M 39.8M 34.2M 45.1M 47.9M 38.4M 249.5M
CPT 5.0% 169.4M 227.8M 119.8M 232.7M 235.5M 226.1M 1.2B
SFT - 18.8K 39.1K 1.6K 5.3K 5.9K 7.5K 78.2K
A.2 Detailed Setup on MMedBench
Data for Continual Pre-Training. To investigate high-quality domain knowledge injection for LLMs and the scalability of injection methods, we curate a series of training corpus sizes from the 25.5B MMedC (Qiu et al., 2024) dataset, including 0.1%, 0.3%, 0.5%, 1%, and 5%, which respectively takes 30M, 76M, 132M, 250M, and 1.2B tokens. We sorted the document-level text content (including textbooks and other corpus from websites and wikipedia) by length in descending order and progressively included more samples to expand the training dataset at larger scales, where Tab. A3 provides detailed statistics. In particular, as MMedC does not provide English textbooks in its released data (due to copyright issues), we collect 18 English textbooks (Jin et al., 2020) as an alternative for English medical knowledge injection, since they have a common OCR source with MMedC Qiu et al. (2024). These English textbooks take around 21.5M tokens.
Knowledge Structure Extraction. For textbooks, we split the data to chunks (knowledge points) within 3072 tokens, and use our specifically developed 7B-size LLM to extract the structured medical knowledge system. For non-textbook data (for instance, MMedC does not provide Japanese textbooks), a clustering-based technique (Sarthi et al., 2024) is adopted to recursively build knowledge structures from fragmented text segments. Fig. A1 presents an example to illustrate the two kinds of knowledge structure extraction processes, Tab. A4 displays the overall statistics on extracted knowledge structures for the final 1.2B tokens.
Data for Supervised Fine-Tuning. As introduced in Sec. 3.3, we prompt Llama3-70B (Dubey et al., 2024) to build the structure-aware answer explanations on top of the raw SFT samples in MMedBench’s training split, and generate extra QA pairs by traversing the extracted knowledge structure from textbooks. The final quantity statistics are presented in Tab. A3. Note that the 70B-size model is not necessary to synthesize QAs and explanations, since we have provided relevant text sources (retrieved from the training corpus) to supplement medical knowledge. In this way, LLMs only need to perform in-context comprehension, rather than generate new QAs based on their own knowledge, which can also be achieved by smaller models like Qwen2.5-7B Yang et al. (2024).
Table A4: Knowledge structure on 1.2B training corpus.
Lang. Book Chapter Section KnowledgePoints
6 2933 14,411 23,239 180,793
A.3 Terminology Explanation
Figure A2: Definition and example of knowledge paths and branches.
Knowledge Paths and Branches. Fig. A2 shows an example of how we define the knowledge paths and branches of the extracted knowledge structure for SSFT data synthesis.
  • A path means a knowledge path from the domain summary (e.g., Biochemistry) to specific knowledge points (e.g., Lipid Metabolism and Cholesterol Transport): “Biochemistry – Overview of lipoprotein metabolism, hormone synthesis – Lipoprotein Metabolism – Lipid Metabolism and Cholesterol Transport”.
  • A branch means the knowledge branch of the tree structure. If a question is related to two knowledge points (e.g., Lipid Metabolism and Cholesterol Transport and Pancreatic zymogen activation) at different branches of the knowledge tree, the knowledge path contains two branches, which becomes the right-bottom part of Fig. A2.
A.4 Resource Requirement
Appendix B Additional Experiments
B.1 Knowledge Structure Extraction
Extracting domain knowledge structure is a prerequisite for subsequent knowledge injection (including both SCPT and SSFT) for language models. In Sec. 3.1, we propose a bottom-up strategy to re-chunk the texts from domain textbooks, summarize a title for each chunk, and send the title list to a specialized 7B model to derive the knowledge structure. The prompt template is displayed in Appendix D.
In fact, we argue that due to the language diversity, a perfectly recovered table of contents of textbooks is unnecessary for domain knowledge injection. A reasonable knowledge structure is sufficient enough. In Tab. A5, we individually adopt few-shot GPT-3.5-Turbo (Brown et al., 2020) and LLaMA3-70B (Dubey et al., 2024) models to extract medical knowledge structure from 18 English textbooks (Jin et al., 2020) (with 26.1M tokens) for subsequent knowledge injection (the backbone LLM is LLaMA2-7B (Touvron et al., 2023b)). Although they present 3.6%-3.7% enhancement on MMedBench (Qiu et al., 2024)’s English test set (denoted as “improvement”), leveraging GPT-3.5-Turbo and LLaMA3-70B is either expensive or time-consuming. GPT-3.5-Turbo costs around 15 dollars to process 26M tokens, while LLaMA3-70B takes around 1.5 hours on 2 A100-80G GPUs, of which both limit scaling data pre-processing for structure-aware knowledge injection.
Model Improvement Time Cost Extra Fee
GPT-3.5-Turbo +3.58 - 15\$
LLaMA3-70B +3.72 1.5h -
Ours-7B +3.69 0.2h -